one side of a square is increased by 30% to maintain the same area ,the other side will have to be decreases by
Answer by ankor@dixie-net.com(22740)

one side of a square is increased by 30%; to maintain the same area, the other side will have to be decreases by
:
L & W = original dimensions
then
1.3L = the increase dimension
;
x = amt in decimals multiplied by the other side to maintain area
:
Original area = new area
L * W = 1.3L * xW
W = 1.3 * xW
xW = W/1.3
x = 1/1.3
x ~ .77
(1-.77) * 100 = 23% reduction in the other dimension